OpenStudy (anonymous):

Metal&Fan. Find the exact value of the radical expression in simplest form. the square root of p to the sixth power + the square root of 9 times p to the sixth power + 3 the square root of p to the sixth power − the square root of p squared 13 times p cubed times the square root of p − 2 times p times the square root of p 13 times p cubed times the square root of p − 2p 7p3 7p3 − p

OpenStudy (johnweldon1993):

\[\large \sqrt{9p^6} + 3\sqrt{p^6} - \sqrt{p^2}\] Hint...what is the square root of 9? it's 3 right? so we can take that out \[\large 3\sqrt{p^6} + 3\sqrt{p^6} - \sqrt{p^2}\] We can combine those first 2 radicals now...add the 3's and leave the radicals alone for now \[\large 6\sqrt{p^6} - \sqrt{p^2}\] any idea what to do next?

OpenStudy (anonymous):

Simplify the expressions both inside and outside the radical by multiplying.?

OpenStudy (anonymous):

@midhun.madhu1987